C90  ^{Position: Chrome}^{Date: 31.07.91}  ^{Date: 12.08.91}^{NR: Dolby C}  ^{NR: Dolby C}^{Src: CD}  ^{Src: CD} The Beatles  Magical Mystery Tour Revolver^{Magical Mystery Tour}^{The Fool On The Hill}^{Flying}^{Blue Jay Way}^{Your Mother Should Know}^{I Am The Walrus}^{Hello Goodbye}^{Strawberry Fields Forever}^{Penny Lane}^{Baby You’re A Rich Man}^{All You Need Is Love}  ^{Taxman}^{Eleanor Rigby}^{I’m Only Sleeping}^{Love You Too}^{Here, There And Everywhere}^{Yellow Submarine}^{She Said She Said}^{Good Day Sunshine}^{And Your Bird Can Sing}^{For No One}^{Doctor Robert}^{I Want To Tell You}^{Got To Get You Into My Life}^{Tomorrow Never Knows}